OpenStudy (anonymous):

find x such that the distance between the points is 5. (0,0),(x, -4)

OpenStudy (anonymous):

i applied the distance formula, setting it equal to 5

OpenStudy (amistre64):

or just use a vector

OpenStudy (amistre64):

distance formula is the same either way

OpenStudy (amistre64):

x^2 + 16 = 26 x = 3 or -3

OpenStudy (amistre64):

and by 26 i mean 25

OpenStudy (anonymous):

\[\sqrt{{x^2} + 4^2} = 5\]

OpenStudy (amistre64):

right; now square both sides; or simply forgo that steps by doing it in your mind :)

OpenStudy (anonymous):

right so we get: x = +-3

OpenStudy (anonymous):

I thought if i applied the distance formula again with x = 3, i should get 5

OpenStudy (amistre64):

you will

OpenStudy (amistre64):

9+16 = 5^2 25 = 25

OpenStudy (anonymous):

right i see, thank you!